1. A budget for the State of Wyoming has yet to be approved, but there have been many proposals. Below is the voting breakdown on the latest proposal. Determine if there is an association between party affiliation and vote toward the latest budget package. Assume a-0.05. Report all answers to two decimal places if the answer involves decimals.

Favor Budget Proposal Oppose Budget Proposal
Democrats 22 45
Republicans 18 10
Answer Box
Null Hypothesis:   
X2observed =
X2critical =
degrees of freedom =
Accept or Reject Ho
APA-style results sentence:
Answer #1

Null hypothesis: party affiliation and vote toward the latest budget package are independent

Applying chi square test:

Expected Ei=row total*column total/grand total Favour Oppose Total
Democrats 28.21 38.79 67
Republicans 11.79 16.21 28
total 40 55 95
chi square    χ2 =(Oi-Ei)2/Ei Favour Oppose Total
Democrats 1.367 0.994 2.362
Republicans 3.272 2.379 5.651
total 4.639 3.374 8.0126

X2 obsereved =8.01

X2critical =3.841

degree of freedom(df) =(rows-1)*(columns-1)= 1

reject Ho

APA-style results : p value <0.05 , conclude party affiliation and vote toward the latest budget package are related